OpenStudy (anonymous):

evaluate the integral from 2 to 3: 1/(x^(2)-1)

OpenStudy (lalaly):

\[x^2-1= (x-1)(x+1)\]\[\int\limits{\frac{dx}{x^2-1}}\]\[=\int\limits{\frac{dx}{(x-1)(x+1)}}\] use paritial fractions to break \[\frac{1}{(x-1)(x+1)}\] then integrate

myininaya (myininaya):

\[\frac{1}{(x-1)(x+1)}=\frac{A}{x-1}+\frac{B}{x+1}=\frac{A(x+1)+B(x-1)}{(x-1)(x+1)}\] so this means we have \[1=A(x+1)+B(x-1)=Ax+Bx+A-B=(A+B)x+(A-B)\] => \[A+B=0\] \[A-B=1\]

OpenStudy (anonymous):

you can also use a hyperbolic substitution x = cosh theta

myininaya (myininaya):

\[A-B=1=> A=1+B\] \[A+B=0=> (1+B)+B=0=> 1+2B=0 => B=\frac{-1}{2}\] \[A=1+B=> A=1+\frac{-1}{2}=\frac{1}{2}\]

OpenStudy (anonymous):

A+B=0 A−B=1 How do u know this?

myininaya (myininaya):

remember left side=right side => coefficient of x on left side=coefficient of x on right side constant term on left side=constant on right side \[\int\limits_{}^{}(\frac{1}{2} \cdot \frac{1}{x-1}+\frac{-1}{2} \cdot \frac{1}{x+1}) dx\]

myininaya (myininaya):

0x+1=(A+B)x+(A-B) how can this statement be true?

myininaya (myininaya):

if what?

myininaya (myininaya):

the coefficient of x on left side must equal coefficient of x on right side the constant term on left must equal constant term on right

myininaya (myininaya):

we want theses expressions to be exactly the same 0x+1 and (A+B)x+(A-B)

myininaya (myininaya):

the only way this can happen is if A+B=0 and if A-B=1

myininaya (myininaya):

do you understand? i don't think i can be anymore clear i'm sorry

OpenStudy (anonymous):

lol yes, sorry it just took me a minute but that makes perfect sense, i see it now....so now that i have my a and b...i plug it back in, right

myininaya (myininaya):

yes

myininaya (myininaya):

\[\int\limits\limits_{}^{}(\frac{1}{2} \cdot \frac{1}{x-1}+\frac{-1}{2} \cdot \frac{1}{x+1}) dx \]

myininaya (myininaya):

\[\frac{1}{2} \ln|x-1| -\frac{1}{2} \ln|x+1|+C\]

myininaya (myininaya):

ok but we started with a definite integral

OpenStudy (anonymous):

correct

myininaya (myininaya):

\[\int\limits_{2}^{3}\frac{1}{x^2-1} dx=[\frac{1}{2}\ln|x-1|-\frac{1}{2}\ln|x+1|]_2^3\]

myininaya (myininaya):

\[=(\frac{1}{2}\ln|3-1|-\frac{1}{2} \ln|3+1|)-(\frac{1}{2}\ln|2-1|-\frac{1}{2} \ln|2+1|)\]

myininaya (myininaya):

\[=\frac{1}{2}\ln|2|-\frac{1}{2}\ln|4|-\frac{1}{2}\ln|1|+\frac{1}{2}\ln|3|\]

myininaya (myininaya):

\[=\frac{1}{2}\ln(2)-\frac{1}{2}\ln(4)-0+\frac{1}{2}\ln(3)\]

myininaya (myininaya):

\[=\ln(\sqrt{2})-\ln(\sqrt{4})+\ln(\sqrt{3})=\ln(\sqrt{2})-\ln(2)+\ln(\sqrt{3})=\ln(\frac{\sqrt{2} \sqrt{3}}{2})\] \[=\ln(\frac{\sqrt{6}}{2})\]

OpenStudy (anonymous):

thank u! I understand this one now haha

myininaya (myininaya):

i'm happy for you :)

OpenStudy (akshay_budhkar):

i wish i could fan myinanaya twice and give her medal 5 times for every question :)

OpenStudy (anonymous):

method using x = cosh theta: let x = cosh theta dx/d theta = sinh theta \[\frac{1}{x^2-1} = \frac{1}{\cosh^2 \theta - 1}=\frac{1}{\sinh^2 \theta}\] Therefore \begin{eqnarray*} \int{\frac{1}{x^2-1}}dx&=&\int{\frac{1}{\sinh^2 \theta}}\sinh \theta d\theta \\ &=&\int{\cosech \theta}d\theta \\ &=&-\coth \theta\cosech \theta + C \end{eqnarray*} Since \[\theta = \cosh^{-1} x\] answer is \[\coth (\cosh^{-1} x) \cosech (\cosh^{-1} x) + C\]

myininaya (myininaya):

eww hyperbolic trig functions

OpenStudy (anonymous):

forgot the minus sign at the very end

myininaya (myininaya):

good job xactxx

OpenStudy (anonymous):

result should be the same since inverse hyperbolic functions are in terms of ln as well ;)

myininaya (myininaya):

yes i would check but i don't want too lol

myininaya (myininaya):

i believe you

OpenStudy (anonymous):

let me see \[-\coth(\cosh^{−1}3)\cosech(\cosh^{−1}3) +\coth(\cosh^{−1}2)\cosech(\cosh^{−1}2)\] cosh^-1 3 = 1.763, cosh^-1 2 = 1.317. so \[-\coth 1.763 \cosech 1.763 + \coth 1.317 \cosech 1.317\] \[=-1.061 x 0.353 + 1.732 x 0.577\] \[=-0.665 + 1\] \[=0.334\] Possibly correct answer is 1/3?

myininaya (myininaya):

they are saying the antiderivative is negative tan hyperbolic inverse+C